A rental property has an annual income of $26,400 and annual expenses of $7,200. If the property can be sold for $255,000 at the end of the 12 years and i=8%, how much should you be willing to pay for it now?
Answer #1

annual cash flow is=26400-7200=19200

how much should you be willing to pay for it now

=P*((1-(1+r)^(-n))/r)+FV/(1+r)^n

=19200*((1-(1+8%)^(-12))/8%)+255000/(1+8%)^12

=245956.71

the above is answer..